# Stability of Non-Hyperbolic Sets

### v1

@@ 4.2 The secondary tangencies
    $m_{5}=(m_{5,x},m_{5,y})\in W^{u}_{\text{loc}}(z^{(2)})$, $F_{t,a}^{-N}(m_{6})=m_{5}$. By Lemma 7, Remark 9 and Remark 10,
    $(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right)$ $\leq$
    $|m_{5,y}-c_{n,y}|$ $\leq$ (4.20)
    $\frac{T}{\mu^{n_{0}/2}}+O\left(\mu^{n}\lambda^{\theta n}\right),$
    and by (4.13)
    $|m_{5,x}-z^{(2)}_{x}|=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.21)
    In particular, when $a=a_{n}(t)$,
    $|m_{5,y}-c_{n,y}|=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\mu^{n}\lambda^{2\theta n}\right).$ (4.22)
    $m_{4}=(m_{4,x},m_{4,y})\in W^{u}_{\text{loc}}(z^{(1)})$, $F_{t,a}^{-n}(m_{5})=m_{4}$. By Lemma 6 and the fact that $z_{n,y}$ can be estimated by $\mu^{-n}$, see Remark 9,
    $|m_{4,y}|=O\left(\frac{1}{\mu^{n}}\right),$ (4.23)
    and by (4.21), using the linear map backward and $(F5)$ (which ensures that the vector $m_{4,x}-z^{(1)}_{x}$ has a dominating first component) we get,
    $|m_{4,x}-z^{(1)}_{x}|=O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.24)
    $m_{3}=(m_{3,x},m_{3,y})\in W^{u}_{\text{loc}}(c^{\prime})$, $F_{t,a}^{-N}(m_{4})=m_{3}$. By Lemma 6 (see Figure 3) and (4.24),
    $|m_{3}-c|$ $=$ $O\left({\lambda^{\theta{n}}}\right),$ (4.25)
    $|m_{3}-c^{\prime}|$ $=$ $O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.26)
@@
    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}}}\leq|m_{2,y}|\leq K_{2}\frac{1}{\mu^{\theta{n}}}.$ (4.27)
    $m_{1}=(m_{1,x},m_{1,y})\in W^{u}_{\text{loc}}(c)$, $F_{t,a}^{-N}(m_{2})=m_{1}$. By (4.27) and because of the quadratic behavior, there exists a uniform constant $K_{1}>0$ such that
    $\frac{1}{K_{1}}\frac{1}{\mu^{\theta\frac{n}{2}}}\leq|m_{1,y}-c_{y}|\leq K_{1}\frac{1}{\mu^{\theta\frac{n}{2}}}.$ (4.28)
Let us recall that, $m_{2}=F^{N}(m_{1})$, $m_{3}=F^{\theta n}(m_{2})$, $m_{4}=F^{N}(m_{3})$, $m_{5}=F^{n}(m_{4})$, and $m_{6}=F^{N}(m_{5})$. Moreover, $t_{1}\in[-t_{0},t_{0}]$ was chosen arbitrary. Take $(t_{1}+t,a)\in\mathcal{B}_{n}$ and consider the ${\mathcal{C}}^{4}$} map
@@
$dm_{i+1}$ $=$ $DF^{N}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.29)
@@
$dm_{i+1}$ $=$ $DF^{n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.30)
@@
$dm_{i+1}$ $=$ $DF^{\theta n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial a}da.$ (4.31)
All partial derivatives in (4.29) are uniformly bounded. However we will need more careful estimate for $\Delta m_{i+1,y}$. Namely, for $i=1,3,5$,
$DF^{N}_{t,a}\left(m_{i}\right)$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(|m_{i,x}|\right)+D_{i}(m_{i,y}-1)\end{matrix}\right),$
which follows from (4.19). Similarly from (4.19) one obtains
@@
$=$ $d_{i}\left|m_{i,y}-1\right|^{2}+O\left(|m_{i,x}|\right),$
where $d_{i}\neq 0$ is uniformly bounded away from zero. By differentiation of (4.19) with respect of $\Delta a$ we obtain
$\frac{\partial m_{i+1,y}}{\partial a}$ $=$ $1+O\left(\left|m_{i,y}-1\right|^{2}\right)+O\left(|m_{i,x}|\right).$
@@
$DF^{N}_{t,a}\left(m_{3}\right)$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(|m_{3}-c^{\prime}|\right)\end{matrix}\right),$
@@
$dm_{i+1}$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(|m_{i,x}|\right)+D_{i}(m_{i,y}-1)\end{matrix}\right)dm_{i}$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{i}|m_{i,y}-1|^{2}+O\left(|m_{i,x}|\right)\right]dt+\left[1+O\left(|m_{i,y}-1|^{2}+|m_{i,x}|\right)\right]da\end{matrix}\right),$
@@
$dm_{4}$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(|m_{3}-c^{\prime}|\right)\end{matrix}\right)dm_{3}$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{3}|m_{3,y}-1|^{2}+O\left(|m_{3,x}|\right)\right]dt+\left[1+O\left(|m_{3,y}-1|^{2}+|m_{3,x}|\right)\right]da\end{matrix}\right),$
@@
$C_{3}=C(1+\lambda^{\theta_{n}}).$ (4.34)
From $dm_{4}$ to $dm_{5}$ we use the linear map $F^{n}$. From (4.30) and using that $m_{5,y}=\mu^{n}m_{4,y}$ we get
$dm_{5}$ $=$ $\left(\begin{matrix}\lambda^{n}&0\\ 0&\mu^{n}\end{matrix}\right)dm_{4}+\left(\begin{matrix}X_{4}\frac{n\lambda^{n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{5,y}\frac{n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$ (4.35)
with $0\neq X_{4}=m_{4,x}\approx 1$. A similar formula holds going from $dm_{2}$ to $dm_{3}$ where we use the linear map $F^{\theta n}$, (4.31) and the fact that $m_{3,y}=\mu^{\theta n}m_{2,y}$. Namely,
$dm_{3}$ $=$ $\left(\begin{matrix}\lambda^{\theta n}&0\\ 0&\mu^{\theta n}\end{matrix}\right)dm_{2}+\left(\begin{matrix}X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{3,y}\frac{\theta n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right)$ (4.36)
@@
$0\neq X_{2}=m_{2,x}\approx\left(\begin{matrix}1&0&\ldots&0\end{matrix}\right)$ (4.37)
Observe that $m_{1,x}=0$ and recall that $\left(dm_{1,x},dm_{1,y}\right)=\left(0,dy\right)$. By (4.2) and by (4.28),
@@
where $\tilde{K}_{1}$ is the constant which gives an equality in (4.28). As a consequence
@@
where $\tilde{d}_{1}=d_{1}\tilde{K}_{1}^{2}$ and $\tilde{D}_{1}=D_{1}\tilde{K}_{1}$. By using the fact that $F^{\theta n}_{t,a}$ is linear and using (4.36),
@@
where $X_{2}\neq 0$ is pointing in the direction of $e_{1}$, see Figure 3. By (4.2) and (4.25) (observe that $\left\{|m_{3,y}-c_{n,y}|,|m_{3,x}|\right\}=O\left(\lambda^{\theta n}\right)$), we get
@@
As consequence, using also (4.37) and (4.34),
@@
$+$ $O\left(\lambda|^{\theta{n}}dy\right).$
By (4.35) we get
@@
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
where we used that $m_{5,y}=(m_{5,y}-c_{n,y})+c_{n,y}$, the fact that the point $c$ is at height 11 and (-4.2). Similarly,
$dm_{5,y}$ $\leq$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
$+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right).$
In particular, when $a=a_{n}$, using the same equalities as before and 4.22,
@@
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
By (4.2), the lower bound in (-4.2) and (4.21) we get
$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\geq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&O\left(\lambda^{n}\right)+D_{5}(C-E)\left(\lambda|^{\theta}\mu\right)^{{n}}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{5}(C-E)^{2}\left(\lambda|^{\theta}\mu\right)^{{2n}}+O\left(\lambda^{n}\right)\right]dt+\left[1+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)\right]da\end{matrix}\right).$
As consequence,
@@
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
$dm_{6,y}$ $\geq$ $\left[D_{5}(C-E)\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}(C-E)n\left(\lambda^{\theta}\mu\right)^{2n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]dt$
$+$ $\left[D_{5}(C-E)\left(\lambda^{\theta}\mu^{2}\right)^{n}+1+D_{5}(C-E)\left(\lambda^{\theta}\mu\right)^{n}\frac{n}{\mu}\frac{\partial\mu}{\partial a}\right$.
$+$ $\left.D_{5}(C-E)n\left(\lambda^{\theta}\mu\right)^{2n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
$+$ $\left.O\left(\left(\lambda^{\theta}\mu\right)^{2n}\right)\right]da+O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
where $D_{5},B_{5},(C-E)\neq 0$ and $\lambda^{\theta}\mu^{2}>1$ (see (2.16)). Similarly, By (4.2), the upper bound in (-4.2) and (4.21) we get
$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\leq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&{D_{5}T}/{\mu^{n_{0}/2}}+O\left(\lambda^{\theta}\mu\right)^{n}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[{d_{5}T^{2}}/{\mu^{n_{0}}}+O\left(\lambda^{\theta}\mu\right)^{2n}\right]dt+\left[1+O\left({1}/{\mu^{n_{0}}}\right)\right]da\end{matrix}\right)$
@@
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$
$dm_{6,y}$ $\leq$ $\left[\frac{D_{5}T}{\mu^{n_{0}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{D_{5}T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n_{0}}}\right)\right]dt$ (4.43)
$+$ $\left[D_{5}T\mu^{n-n_{0}/2}+1+\frac{D_{5}T}{\mu^{n_{0}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{D_{5}T^{2}}{\mu^{n_{0}}}+O\left(\frac{1}{\mu^{n_{0}}}\right)\right]da$
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{2n}}\right)dy,$
where $D_{5},B_{5},T\neq 0$ and $\mu^{n-n_{0}/2}>1$ (see (2.16)).
In particular, when $a=a_{n}$, using the same equalities as before, (4.22) and (4.2) we get,
@@
$+$ $O\left(\left(\lambda|^{\theta}\mu\right)^{{n}}dy\right),$

### v3

@@ 4.2 The secondary tangencies
    $m_{5}=(m_{5,x},m_{5,y})\in W^{u}_{\text{loc}}(z^{(2)})$, $F_{t,a}^{-N}(m_{6})=m_{5}$. By Lemma 4.1.21 and (4.0.1),
    $(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$ $\leq$
    $m_{5,y}-c_{n,y}$ $\leq$ (4.2.4)
    $\frac{T}{\mu^{n_{0}^{*}/2}}+O\left(\frac{n}{\mu^{n+{n_{0}^{*}}/{2}}}\right),$
    and by (4.1.31)
    $m_{5,x}-z^{(2)}_{x}=O\left(\left(\frac{\lambda^{2-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.2.5)
    In particular, when $a=a_{n}(t)$ by Lemma 4.1.21 and (4.1.32),
    $m_{5,y}-c_{n,y}=C\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right).$ (4.2.6)
    $m_{4}=(m_{4,x},m_{4,y})\in W^{u}_{\text{loc}}(z^{(1)})$, $F_{t,a}^{-n}(m_{5})=m_{4}$. By Lemma 4.1.10 and the fact that $z_{n,y}$ can be estimated by $\mu^{-n}$, see Remark 4.0.3,
    $m_{4,y}=O\left(\frac{1}{\mu^{n}}\right),$ (4.2.7)
    and by (4.2.5), using the linear map backward we get,
    $m_{4,x}-z^{(1)}_{x}=O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.2.8)
    $m_{3}=(m_{3,x},m_{3,y})\in W^{u}_{\text{loc}}(c^{\prime})$, $F_{t,a}^{-N}(m_{4})=m_{3}$. By (4.1.7) and (4.2.8),
    $m_{3}-c$ $=$ $O\left({\lambda^{\theta{n}}}\right),$ (4.2.9)
    $m_{3}-c^{\prime}$ $=$ $O\left(\left(\frac{\lambda^{1-\theta}}{\mu^{2}}\right)^{n}\right).$ (4.2.10)
@@
    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}}}\leq m_{2,y}\leq K_{2}\frac{1}{\mu^{\theta{n}}}.$ (4.2.11)
    $m_{1}=(m_{1,x},m_{1,y})\in W^{u}_{\text{loc}}(c)$, $F_{t,a}^{-N}(m_{2})=m_{1}$. By (4.2.11), (2.2.10) and Remark 4.0.3,
    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}}}\leq m_{2,y}-z_{y}\leq K_{2}\frac{1}{\mu^{\theta{n}}}.$
    and because of the quadratic behavior, see Lemma 4.1.15,
    $\frac{1}{K_{2}}\frac{1}{\mu^{\theta{n}/2}}\leq m_{2,x}-1\leq K_{2}\frac{1}{\mu^{\theta{n}/2}}$
    where we adjusted the value of the constant $K_{2}$. Hence, there exists a uniform constant $K_{1}>0$ such that
    $\frac{1}{K_{1}}\frac{1}{\mu^{\theta\frac{n}{2}}}\leq m_{1,y}-c_{y}\leq K_{1}\frac{1}{\mu^{\theta\frac{n}{2}}}.$ (4.2.12)
Let us recall that, $m_{2}=F^{N}(m_{1})$, $m_{3}=F^{\theta n}(m_{2})$, $m_{4}=F^{N}(m_{3})$, $m_{5}=F^{n}(m_{4})$, and $m_{6}=F^{N}(m_{5})$. Moreover, $t_{1}\in[-t_{0},t_{0}]$ was chosen arbitrarly. Take $(t_{1}+\Delta t,a)\in\mathcal{B}_{n}$ and consider the ${\mathcal{C}}^{4}$} map
@@
$dm_{i+1}$ $=$ $DF^{N}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{N}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.2.13)
@@
$dm_{i+1}$ $=$ $DF^{n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{n}_{t,a}\left(m_{i}\right)}{\partial a}da,$ (4.2.14)
@@
$dm_{i+1}$ $=$ $DF^{\theta n}_{t,a}\left(m_{i}\right)dm_{i}+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial t}dt+\frac{\partial F^{\theta n}_{t,a}\left(m_{i}\right)}{\partial a}da.$ (4.2.15)
All partial derivatives in (4.2.13) are uniformly bounded. However we will need more careful estimate for $\Delta m_{i+1,y}$. Namely, for $i=1,3,5$,
$DF^{N}_{t,a}\left(m_{i}\right)$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(m_{i,x}\right)+D_{i}(m_{i,y}-1)\end{matrix}\right),$
which follows from (4.2.3). Similarly from (4.2.3) one obtains
@@
$=$ $d_{i}\left(m_{i,y}-1\right)^{2}+O\left(m_{i,x}\right),$
where $d_{i}\neq 0$ is uniformly bounded away from zero. By differentiation of (4.2.3) with respect to $\Delta a$ we obtain
$\frac{\partial m_{i+1,y}}{\partial a}$ $=$ $1+O\left(\left(m_{i,y}-1\right)^{2}\right)+O\left(m_{i,x}\right).$
@@
$DF^{N}_{t,a}\left(m_{3}\right)$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(m_{3}-c^{\prime}\right)\end{matrix}\right),$
@@
$dm_{i+1}$ $=$ $\left(\begin{matrix}A_{i}&B_{i}\\ C_{i}&O\left(m_{i,x}\right)+D_{i}(m_{i,y}-1)\end{matrix}\right)dm_{i}$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{i}(m_{i,y}-1)^{2}+O\left(m_{i,x}\right)\right]dt+\left[1+O\left((m_{i,y}-1)^{2}+m_{i,x}\right)\right]da\end{matrix}\right),$
@@
$dm_{4}$ $=$ $\left(\begin{matrix}A_{3}&B_{3}\\ C_{3}&O\left(m_{3}-c^{\prime}\right)\end{matrix}\right)dm_{3}$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{3}(m_{3,y}-1)^{2}+O\left(m_{3,x}\right)\right]dt+\left[1+O\left((m_{3,y}-1)^{2}+m_{3,x}\right)\right]da\end{matrix}\right),$
@@
$C_{3}=C(1+O\left(\lambda^{\theta_{n}}\right)).$ (4.2.18)
From $dm_{4}$ to $dm_{5}$ we use the linear map $F^{n}$. From (4.2.14) and using that $m_{5,y}=\mu^{n}m_{4,y}$ we get
$dm_{5}$ $=$ $\left(\begin{matrix}\lambda^{n}&0\\ 0&\mu^{n}\end{matrix}\right)dm_{4}+\left(\begin{matrix}X_{4}\frac{n\lambda^{n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{5,y}\frac{n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$ (4.2.19)
with $0\neq X_{4}=m_{4,x}\approx 1$. A similar formula holds going from $dm_{2}$ to $dm_{3}$ where we use the linear map $F^{\theta n}$, (4.2.15) and the fact that $m_{3,y}=\mu^{\theta n}m_{2,y}$. Namely,
$dm_{3}$ $=$ $\left(\begin{matrix}\lambda^{\theta n}&0\\ 0&\mu^{\theta n}\end{matrix}\right)dm_{2}+\left(\begin{matrix}X_{2}\frac{\theta n\lambda^{\theta n}}{\lambda}\left[\frac{\partial\lambda}{\partial t}dt+\frac{\partial\lambda}{\partial a}da\right]\\ m_{3,y}\frac{\theta n}{\mu}\left[\frac{\partial\mu}{\partial t}dt+\frac{\partial\mu}{\partial a}da\right]\end{matrix}\right),$ (4.2.20)
@@
$0\neq X_{2}=m_{2,x}\approx 1.$ (4.2.21)
Observe that $m_{1,x}=0$ and recall that $\left(dm_{1,x},dm_{1,y}\right)=\left(0,dy\right)$. By (4.2) and by (4.2.12),
@@
where $\tilde{K}_{1}$ is the constant which gives an equality in (4.2.12). As a consequence
@@
where $\tilde{d}_{1}=d_{1}\tilde{K}_{1}^{2}$ and $\tilde{D}_{1}=D_{1}\tilde{K}_{1}$. By using the fact that $F^{\theta n}_{t,a}$ is linear and using (4.2.20),
@@
where $X_{2}\neq 0$. By (4.2) and (4.2.9) (observe that $m_{3,y}-c_{n,y},m_{3,x}=O\left(\lambda^{\theta n}\right)$), we get
@@
As consequence, using also (4.2.21) and (4.2.18),
@@
$+$ $O\left(\lambda^{\theta{n}}dy\right).$
By (4.2.19) we get
@@
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
where we used that $m_{5,y}=(m_{5,y}-c_{n,y})+c_{n,y}$, the fact that the point $c$ is at height $1$ and (-4.2). Similarly,
$dm_{5,y}$ $\leq$ $\left[\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{T}{\mu^{n_{0}*/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
$+$ $\left[\mu^{n}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{T}{\mu^{n_{0}^{*}/2}}+n\left(\lambda^{\theta}\mu\right)^{n}C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+O\left(\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right).$
In particular, when $a=a_{n}$, using the same equalities as before and 4.2.6,
@@
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
By (4.2), the lower bound in (-4.2) and (4.2.5) we get
$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\geq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&D_{5}S_{C-E}\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[d_{5}S_{C-E}^{2}\right]dt+\left[1+O\left(S_{C-E}^{2}\right)\right]da\end{matrix}\right),$
where $S_{C-E}=(C-E)\left(\lambda^{\theta}\mu\right)^{n}+O\left(\left(\lambda^{\theta}\mu^{1-\frac{\theta}{2}}\right)^{n}\right)$. As consequence,
@@
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
$dm_{6,y}$ $\geq$ $\left[D_{5}S_{C-E}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+D_{5}S_{C-E}n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial t}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial t}\right]\right$.
$+$ $\left.O\left(S_{C-E}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]dt$
$+$ $\left[D_{5}S_{C-E}\mu^{n}+1+D_{5}S_{C-E}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+D_{5}S_{C-E}n\left(\lambda^{\theta}\mu\right)^{n}\left[C\frac{\theta}{\lambda}\frac{\partial\lambda}{\partial a}+\frac{(C-E)}{\mu}\frac{\partial\mu}{\partial a}\right]\right$.
$+$ $\left.O\left(S_{C-E}\left(\lambda^{\theta}\mu\right)^{n}\right)\right]da+O\left(\left(\lambda^{\theta}\mu\right)^{{2n}}\right)dy$ (4.2.25)
where $D_{5},B_{5},(C-E)\neq 0$ and $S_{C-E}\mu^{n}>1$ (see (2.3.1)). Similarly, By (4.2), the upper bound in (-4.2) and (4.2.5) we get
$\left(\begin{matrix}dm_{6,x}\\ dm_{6,y}\end{matrix}\right)$ $\leq$ $\left(\begin{matrix}A_{5}&B_{5}\\ C_{5}&{D_{5}T}/{\mu^{n_{0}^{*}/2}}+O\left({n}/{\mu^{n+n_{0}^{*}/2}}\right)\end{matrix}\right)\left(\begin{matrix}dm_{5,x}\\ dm_{5,y}\end{matrix}\right)$
$+$ $\left(\begin{matrix}O\left(dt\right)+O\left(da\right)\\ \left[{d_{5}T^{2}}/{\mu^{n_{0}^{*}}}+O\left({n}/{\mu^{n+n_{0}^{*}}}\right)\right]dt+\left[1+O\left({1}/{\mu^{n_{0}^{*}}}\right)\right]da\end{matrix}\right)$
@@
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$
$dm_{6,y}$ $\leq$ $\left[\frac{D_{5}T}{\mu^{n_{0}^{*}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial t}+\frac{n}{\mu}\frac{\partial\mu}{\partial t}\frac{D_{5}T^{2}}{\mu^{n_{0}^{*}}}+O\left(\frac{1}{\mu^{n_{0}^{*}}}\right)\right]dt$ (4.2.27)
$+$ $\left[D_{5}T\mu^{n-n_{0}^{*}/2}+1+\frac{D_{5}T}{\mu^{n_{0}^{*}/2}}\frac{n}{\mu}\frac{\partial\mu}{\partial a}+\frac{n}{\mu}\frac{\partial\mu}{\partial a}\frac{D_{5}T^{2}}{\mu^{n_{0}^{*}}}+O\left(\frac{1}{\mu^{n_{0}^{*}}}\right)\right]da$
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{2n}}\right)dy,$
where $D_{5},B_{5},T\neq 0$ and $\mu^{n-n_{0}^{*}/2}>1$ (see (4.0.1) and (2.3.1)).
In particular, when $a=a_{n}$, using the same equalities as before, (4.2.6) and (4.2) we get,
@@
$+$ $O\left(\left(\lambda^{\theta}\mu\right)^{{n}}dy\right),$